Two charges each with a magnitude of 1.1⋅10−7 C experience a force of 4.2⋅10−4 N. How far apart are the charges from each other? 


1
Expert's answer
2021-05-26T10:32:38-0400

"F=kq^2\/d^2"


"d=q\sqrt{\frac{k}{F}}=1.1\cdot10^{-7}\sqrt{\frac{9\cdot10^9}{4.2\cdot10^{-4}}}=0.5\ m"